Interactive LearningWare 4.4 provides a review of
Answer #1

a)

acceleration = a = Force /mass = F/m = (7.54 x 106)/(3.22 x 105) = 23.42

b)

direction = 57.7 degree above horizontal

since acceleration is in same direction as the force